Grade 11IIT JEE Entrance Exam

The wavelength corresponding to maximum energy for hydrogen is 91.2 nm. Find the corresponding wavelength for He^+ ion.

dee
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
We know that , 
1/x = R*Z²*[1/(n1)²-1/(n2)²] where x is wavelength.
According to the question , 
1/91.2 = R*1²*[1/(n1)²-1/(n2)²]                        for hydrogen      ..................1
1/x = R*2²*[1/(n1)²-1/(n2)²]                            for helium           ..................2
Dividing the equations 1 and 2, we get,
x/91.2 = 1/4
x = 22.8 nm